# Optimization via Quantum Preconditioning

### v1

@@ IV.1 Quantum-Inspired Advantage for Random $3$-Regular Graph Maximum-Cut Problems
Figure 5: Average time for preconditioning $N$-variable random $3$-regular graph maximum-cut problems using classical state vector emulations of the QAOA using the light-cone technique at $p=1$ and $p=2$. The maximum preconditioning time budget for an advantage via SA for an approximation ratio of $\alpha=99.9\%$ is computed from Fig. 1. Run-times corresponds to a 64Gb MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
@@
We refer to a quantum-inspired advantage as obtained via a purely classical calculation or emulation of a quantum circuit, as opposed to execution on an actual quantum computer. This can be a first step towards genuine quantum utility if the proposed method has room to perform even better at a larger scale (more qubits, larger depth) beyond classical computational capabilities.
@@
We focus on the random $3$-regular graph maximum-cut problems, which present the largest advantage window compared to the SK spin glasses or the grid energy problem. We consider SA, which is more efficient than the BM solver in the $\alpha=99.9\%$ regime (this value has been chosen arbitrarily and others may provided a larger advantage window). The preconditioning time budget is defined in Fig. 1b, as the difference in SA run-time between the original and preconditioned problems to get to $\alpha=99.9\%$ We report the budget between original and preconditioned at $p=2$ in Fig. 5. Because there is about one order of magnitude difference in runtime between the original and preconditioned problems, the budget is mostly dominated by the run-time of the solver on the original problem.
@@
We perform the preconditioning via state vector emulations of the QAOA circuits on the relevant light-cone-induced subgraphs according to Sec. III.2.1. Determining the light-cone induced subgraphs of an $N$-variable problem scales as $O(N)$. The QAOA implementation benefits from the global $\mathbb{Z}_{2}$ sign-flip symmetry of the problem that reduces the overall emulation run-time by a theoretical factor two. Moreover, we use the fact that in the limit $p\ll\ln N$, most light-cone-induced subgraphs are trees (see Appendix E). Trees are $N$-variable graphs for which the number of edges is $n=N-1$. As such, we cache the expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ of the trees such that the QAOA only needs to be emulated once on such subgraphs. This results in only emulating $O(\exp p)$ QAOA circuits, independently of $N$ (see Appendix E). The memory usage and algorithmic complexity of a QAOA emulation via the light cone technique scales as $O(\exp\exp p)$ and classically manageable at $p=1$ and $p=2$.
Figure 6: (A-D) Random $N$-variable $3$-regular graph maximum-cut problems are considered. Convergence of the average approximation ratio $\alpha$ towards its infinite-sampling value, obtained via quantum preconditioning with a finite number of sampled bit strings $K$. Preconditioning is performed via the simulation of light-cone-induced $p$-layer QAOA circuits. (A) Simulated annealing (SA) at $p=1$ with $n_{\textrm{iter}}=100$ iterations. (B) SA at $p=2$ with $n_{\textrm{iter}}=100$ iterations. (C) Burer-Monteiro solver (BM) at $p=1$ with $n_{\textrm{iter}}=50$ iterations. (D) BM solver at $p=2$ with $n_{\textrm{iter}}=50$ iterations. Each data point is averaged over $200$ randomly generated problem instances. (E) Average sampling time on Rigetti Ankaa-3 superconducting quantum chip for a light-cone-induced QAOA circuit for various numbers of QAOA layers $p$ for random $3$-regular graph maximum-cut problems. RESET corresponds to a passive reset time of $200~{}\mu$s of the qubits between two circuit executions. Each data point is averaged over $50$ randomly selected light-cone-induced QAOA circuits. Error bars indicate the standard error of the mean.
@@
We show in Fig. 5 the time it takes to perform the quantum preconditioning at $p=1$ and $p=2$ via emulations. For large enough problems, the preconditioning run-time fits within the budget. Therefore, quantum preconditioning can deliver a quantum-inspired advantage over SA through classical emulations. This advantage extends for larger problem sizes to the BM solver, which has a much larger budget (see Fig. 2)—however, the BM solver is not competitive against SA.
@@
The change of slope observed at large $N$ for $p=1$ is because the run-time becomes dominated by the $O(N)$ task of determining the light-cone induced subgraphs rather than the QAOA emulations. We anticipate an analogous asymptotic behavior for $p=2$.
@@ IV.2 Prospects for Hardware-Based Quantum Utility for Random $3$-Regular Graph Maximum-Cut Problems
We have shown in Fig. 5 that quantum preconditioning can provide a quantum-inspired advantage on random $3$-regular graph maximum-cut problems via classical emulations at $p=1$ and $p=2$. We have also shown in Fig. 1b that increasing the depth of the QAOA from $p=1$ to $p=2$ can accelerate the convergence. Hence, a question is whether $p=3$ could improve the convergence further. However, light-cone induced subgraphs at $p=3$ have up to $43$ qubits and, therefore, cannot be emulated with exact classical methods in a reasonable amount of time, calling for a new simulation platform, such as an actual quantum computer.
@@ IV.2.1 Sampling From the Wave Function
When moving to quantum hardware for executing quantum circuits, one does not have access to the full quantum state $|\Psi\rangle_{p}$ of Eq. (3) for computing expectation values but has to sample bit strings $\{\boldsymbol{z}^{(k)}\}_{k=1,\ldots K}$ to estimate them. After executing the circuit and measuring the qubits $K$ times, one has collected $K$ bit strings according to the probability $|\langle\boldsymbol{z}^{(k)}|\Psi\rangle_{p}|^{2}$. Two-point correlations are then estimated as
@@
The total run-time for estimating the expectation value is proportional to $K$. For the classical state vector emulations, we had access to the full quantum state $|\Psi\rangle_{p}$ and did not need to sample bit strings, which corresponded to the $K\to+\infty$ limit.
@@
A first question is whether quantum preconditioning is robust to finite sampling. We show in Figs 6A- 6D the accuracy of SA and the BM solver on preconditioned $N$-variable random $3$-regular graph maximum-cut problems where preconditioning was performed by sampling $K$ bit strings to estimate correlations via Eq. (8). The correlation matrix was computed via the classical emulation of light-cone induced QAOA subgraphs, as was done previously in Sec. III.2.1. We consider preconditioning at $p=1$ and $p=2$ and the performance of SA and BM with $n_{\textrm{iter}}=100$ and $n_{\textrm{iter}}=50$ iterations, respectively.
We do not observe a clear $N$ or $p$ dependence of the convergence of the average approximation ratio towards its infinite-sampling ($K\to+\infty$) value. However, we can rule out a logarithmic convergence with $K$ (i.e., an exponential number of samples needed to achieve a desired accuracy). Moreover, we note that a convergence to a relative error in average approximation ratio of $0.01\%$ requires at most $K\simeq 10^{4}$ bit strings, which is a typical number used for evaluating expectation values on quantum computers. The saturation observed at $K\sim 10^{4}$ in Figs. 6B and 6C is due to finite statistics in the number of random problem instances considered and for estimating the approximation ratio of the individual problem instances. We expect all the curves to get asymptotically to zero as $K\to+\infty$. Therefore, finite sampling does not seem to be a barrier to the performance of quantum preconditioning.
@@
The next question is the preconditioning time through sampling. We consider the superconducting quantum computer Ankaa-3 developed by Rigetti Computing and accessed via cloud services. Superconducting qubits provide the fastest sampling rate to date thanks to their logical quantum operations in the tens of nanoseconds range, several orders of magnitude faster than atom- or ion-based quantum computers. Given optimization solvers operate on an accuracy versus run-time window, the speed of preconditioning is paramount to deliver any potential advantage.
@@
We report in Fig 6E the average sampling time for the $p$-layer QAOA on a typical $N$-vertex subgraph induced by the light cone on $3$-regular graphs. An arbitrary QAOA circuit on $N$ qubits with $p$ layers executed on a linear chain topology of qubits through a naïve swap network [70] requires a total of $3Np$ compact layers of two-qubit $\texttt{ISWAP}=\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$ gates and $12(2N+1)p+$ 1 compact layers of one-qubit $\texttt{RX}(\phi)=\exp(-i\hat{X}\phi/2)$ gates where $\phi=\pm\pi/2$ [45]. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. All the gates within a layer can be applied in parallel on independent qubits. We also make use of one-qubit $\texttt{RZ}(\theta\in\mathbb{R})=\exp(-i\hat{Z}\theta/2)$ gates, which are virtual and realized by a change of frame of the qubits tracked by the control system. The collection time of $K$ samples is modeled by
$t\bigl{(}K,N,p\bigr{)}=t_{\textrm{ovhead}}\bigl{(}K,N,p\bigr{)}+Kt_{\textrm{circ}}\bigr{(}N,p\bigr{)}$ (9)
@@
where $t_{\textrm{ovhead}}(K,N,p)$ indicates a classical overhead encompassing, e.g., network latency, job management, and control systems. The other component
$t_{\textrm{circ}}\bigl{(}N,p\bigr{)}=3Npt_{\textrm{2Q}}+\bigl{[}2(2N+1)p+1\bigr{]}t_{\textrm{1Q}}+t_{\textrm{mes}}+t_{\textrm{res}}$ (10)
relates to the circuit execution. It defines a minimum physical limit for sampling one quantum circuit. We have $t_{\textrm{2Q}}\simeq 80$ ns the typical duration of a two-qubit gate, $t_{\textrm{1Q}}\simeq 40$ ns the typical duration of a one-qubit gate, $t_{\textrm{mes}}=1~{}\mu$s the typical duration of measuring a qubit, and $t_{\textrm{res}}=200~{}\mu$s the typical duration for reseting passively a qubit prior to a subsequent circuit execution. These numbers provide a baseline but can typically be calibrated or fine-tuned.
@@
For $N=4,096$, the average size of light-cone-induced subgraphs is $N=5.67(2)$ at $p=1$, $N=16.613(5)$ at $p=2$, and $N=39.772(5)$ at $p=3$. Hence, in this regime, the average circuit execution $t_{\textrm{circ}}$ is dominated by the passive reset time of the qubits. It is highlighted in Fig. 6E, where for a large number of samples, $t\simeq Kt_{\textrm{res}}$. For a small number of samples, the sampling time is instead dominated by the classical overhead. For a number of samples $K\gtrsim 10^{3}$, a first step for improving the sampling time is, therefore, to move to an active reset strategy for the qubits, which would yield $t_{\textrm{mes}}+t_{\textrm{res}}\simeq 6~{}\mu$[71], a $30$-fold reduction over passive reset.
@@
At the moment, for $K=10^{4}$, the total sampling time for a single light-cone induced circuit takes about two to three seconds in Fig 6E. This is not competitive against classical state vector emulations at $p=1$ and $p=2$, which can precondition all light-cone-induced circuits in less time—see Fig. 5. Thus, improvements will be needed to reduce the sampling time and make the approach competitive and capable of delivering an advantage when accounting for the preconditioning time budget. Whether this will prove a bottleneck remains to be established: Indeed, we observe in Fig. 5 that the budget increases with the number of variables $N$ while the preconditioning is independent of $N$ for $p\ll\ln N$ thanks to caching of light-cone-induced subcircuits (see Appendix. E)—strictly speaking, preconditioning has a classical $O(N)$ dependence in finding the light-cone induced subgraphs. If confirmed for much larger $N$, this scaling difference between preconditioning budget and time could suggest a crossing value $N^{*}$ for a quantum simulation method (hardware or classically emulated) to deliver an advantage, independently of its intrinsic run-time.
@@ IV.2.2 Performance of a Noisy Quantum Computer
where $F=f^{n_{\textrm{2Q}}}$ can be interpreted as the global quantum circuit fidelity with individual errors happening independently and at random at the two-qubit gate level. In order to distinguish an $O(1)$ expectation value with $K$ sampled bit strings, one needs $K\sim F^{-2}$, or equivalently $K\sim f^{-3pN^{2}}$, according to the central limit theorem. By plugging in relevant numbers for $f$, $N$, and $p$, one realizes that finite sampling by itself is likely not a bottleneck but sampling for reliably estimating noisy and exponentially small expectation values might be. Therefore, hardware performance through $f$ sets the expectation for what circuit can be successfully executed. We verify that such a depolarizing noise model is able to phenomenologically fit experimental data in App. C—also similarly observed in Ref. [55].
@@ IV.2.3 Noisy Quantum Computer Versus Approximate Classical Tensor Networks
Given the fidelity requirements in the absence of quantum error correction to run the QAOA on large problem instances, one will likely need to rely on a decomposition technique such as the light-cone one introduced in Sec. III.2.1. Light-cone-induced subgraphs at $p=3$ require at most $43$ qubits, far beyond what state vector emulators can handle in a reasonable amount of time. Other candidate simulators for simulating such a quantum circuit include an actual quantum computer or classical approximate tensor networks methods, such as matrix product states [72, 73]. In the absence of quantum error correction, quantum computers are imperfect, and thus realize in some aspects, only an approximate simulation. A first question is whether quantum preconditioning via an approximately classically emulated preconditioner can deliver an advantage.
Assuming an imperfect hardware output, a second question is what level of approximation in a tensor network simulation matches the accuracy of the imperfect quantum computer? The level of approximation in tensor networks, encoded in a parameter $\chi\in[1,2^{N/2}]$ controls their run-time: The larger $\chi$ is, the less approximation is performed, and the longer the simulation takes. For instance, the algorithmic complexity of matrix product states goes as $O(\chi^{3})$. Then, one can answer the final question: For a matched output accuracy, which of the approximate tensor network simulator or quantum computer is faster? This will provide a basis for a hardware-based quantum utility [74, 75, 76, 77].
@@
Many challenges faced by the proposed quantum preconditioning method to deliver quantum utility are common to quantum optimization and active areas of research. We discuss some of the main ones in the following.

### v2

@@ IV.1 Quantum-inspired advantage for random $3$-regular graph maximum-cut problems
Figure 5: Average time for preconditioning NN-variable random $3$-regular graph maximum-cut problems using classical state-vector emulations of the QAOA using the light-cone technique at $p=1$ and $p=2$. The maximum preconditioning time budget for an advantage via SA for an approximation ratio of $\alpha=99.9\%$ is computed from Fig. 1. Run-times correspond to a 64GB MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
@@
We refer to a quantum-inspired advantage as obtained via a purely classical calculation or emulation of a quantum circuit, as opposed to execution on an actual quantum computer. This can be a first step toward genuine quantum utility if the proposed method has room to perform even better at a larger scale (more qubits, larger depth) beyond classical computational capabilities.
@@
We focus on the random $3$-regular graph maximum-cut problems, which present the largest advantage window compared to the SK spin glasses or the grid energy problem. We consider SA, which is more efficient than the BM solver in the $\alpha=99.9\%$ regime (this value has been chosen arbitrarily, and others may provide a larger advantage window). The preconditioning time budget is defined in Fig. 1b as the difference in SA run-time between the original and preconditioned problems to reach $\alpha=99.9\%$ . We report the budget between original and preconditioned at $p=2$ in Fig. 5. Because there is a difference of about one order of magnitude in runtime between the original and preconditioned problems, the budget is mostly dominated by the run-time of the solver on the original problem.
@@
We perform the preconditioning via state-vector emulations of the QAOA circuits on the relevant light-cone-induced subgraphs according to Sec. III.1.2. Determining the light-cone-induced subgraphs of an $N$-variable problem scales as $O(N)$. The QAOA implementation benefits from the global $\mathbb{Z}_{2}$ sign-flip symmetry of the problem, which reduces the overall emulation run-time by a theoretical factor two. Moreover, we use the fact that in the limit $p\ll\ln N$, most light-cone-induced subgraphs are trees (see Appendix G). Trees are $N$-variable graphs for which the number of edges is n=N1n=N-1. As such, we cache the expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ of the trees such that the QAOA only needs to be emulated once on such subgraphs. This results in only emulating $O(\exp p)$ QAOA circuits, independent of $N$ (see Appendix G). The memory usage and algorithmic complexity of a QAOA emulation via the light-cone technique scale as $O(\exp\exp p)$ and are classically manageable at $p=1$ and $p=2$.
[…]
@@
In Fig. 5, we show the time it takes to perform the quantum preconditioning at $p=1$ and $p=2$ via emulations. For large enough problems, the preconditioning run-time fits within the budget. Therefore, quantum preconditioning can deliver a quantum-inspired advantage over SA through classical emulations. This advantage extends for larger problem sizes to the BM solver, which has a much larger budget (see Fig. 2); however, the BM solver is not competitive against SA.
@@
The change of slope observed at large $N$ for $p=1$ is because the run-time becomes dominated by the $O(N)$ task of determining the light-cone-induced subgraphs rather than the QAOA emulations. We anticipate an analogous asymptotic behavior for $p=2$.
@@ IV.2 Prospects for hardware-based quantum utility for random $3$-regular graph maximum-cut problems
We have shown in Fig. 5 that quantum preconditioning can provide a quantum-inspired advantage on random $3$-regular graph maximum-cut problems via classical emulations at $p=1$ and $p=2$. We have also shown in Fig. 1b that increasing the depth of the QAOA from $p=1$ to $p=2$ can accelerate the convergence. Hence, a question is whether $p=3$ could improve the convergence further; however, light-cone-induced subgraphs at $p=3$ have up to $43$ qubits and, therefore, cannot be emulated with exact classical methods in a reasonable amount of time, calling for a new simulation platform, such as an actual quantum computer.
@@ IV.2.1 Sampling from the wave function
When moving to quantum hardware for executing quantum circuits, one does not have access to the full quantum state $|\Psi\rangle_{p}$ of Eq. (3) for computing expectation values but must sample bit strings $\{\boldsymbol{z}^{(k)}\}_{k=1,\ldots K}$ to estimate them. After executing the circuit and measuring the qubits $K$ times, one has collected $K$ bit strings according to the probability $|\langle\boldsymbol{z}^{(k)}|\Psi\rangle_{p}|^{2}$. Two-point correlations are then estimated as
@@
The total run-time for estimating the expectation value is proportional to $K$. For the classical state-vector emulations, we had access to the full quantum state $|\Psi\rangle_{p}$ and did not need to sample bit strings, which corresponded to the $K\to+\infty$ limit.
@@
A first question is whether quantum preconditioning is robust to finite sampling. In Figs 6A- 6D, we show the accuracy of SA and the BM solver on preconditioned $N$-variable random $3$-regular graph maximum-cut problems where preconditioning was performed by sampling $K$ bit strings to estimate correlations via Eq. (8). The correlation matrix was computed via the classical emulation of light-cone-induced QAOA subgraphs, as was done previously in Sec. III.1.2. We consider preconditioning at $p=1$ and $p=2$ and the performance of SA and BM with $n_{\textrm{iter}}=100$ and $n_{\textrm{iter}}=50$ iterations, respectively.
We do not observe a clear $N$ or $p$ dependence of the convergence of the average approximation ratio toward its infinite-sampling ($K\to+\infty$) value; however, we can rule out a logarithmic convergence with $K$ (i.e., an exponential number of samples needed to achieve a desired accuracy). Moreover, we note that a convergence to a relative error in average approximation ratio of $0.01\%$ requires at most $K\simeq 10^{4}$ bit strings, which is a typical number used for evaluating expectation values on quantum computers. The saturation observed at $K\sim 10^{4}$ in Figs. 6B and 6C is due to finite statistics in the number of random problem instances considered and for estimating the approximation ratio of the individual problem instances. We expect all the curves to get asymptotically to zero as $K\to+\infty$. Therefore, finite sampling does not seem to be a barrier to the performance of quantum preconditioning.
@@
The next question is the time needed for preconditioning via sampling. We consider the superconducting quantum computer Ankaa-3 developed by Rigetti Computing and accessed via cloud services. Superconducting qubits provide the fastest sampling rate to date thanks to their logical quantum operations in the tens-of-nanoseconds range, several orders of magnitude faster than atom- or ion-based quantum computers. Given that optimization solvers operate on an accuracy versus run-time window, the speed of preconditioning is paramount to deliver any potential advantage.
@@
In Fig 6E, we report the average sampling time for the $p$-layer QAOA on a typical $N$-vertex subgraph induced by the light cone on $3$-regular graphs. An arbitrary QAOA circuit on $N$ qubits with $p$ layers executed on a linear chain topology of qubits through a naïve swap network [69] requires a total of $3Np$ compact layers of two-qubit $\texttt{ISWAP}=\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$ gates and $2(2N+1)p+1$ compact layers of one-qubit $\texttt{RX}(\phi)=\exp(-i\hat{X}\phi/2)$ gates where $\phi=\pm\pi/2$ [44]. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. All the gates within a layer can be applied in parallel on independent qubits. We also make use of one-qubit $\texttt{RZ}(\theta\in\mathbb{R})=\exp(-i\hat{Z}\theta/2)$ gates, which are virtual and realized by a change of frame of the qubits tracked by the control system. The collection time of $K$ samples is modeled by
$t\bigl(K,N,p\bigr)=t_{\textrm{ovhead}}\bigl(K,N,p\bigr)+Kt_{\textrm{circ}}\bigr(N,p\bigr),$ (9)
@@
where $t_{\textrm{ovhead}}(K,N,p)$ indicates a classical overhead encompassing, e.g., network latency, job management, and control systems. The other component,
$t_{\textrm{circ}}\bigl(N,p\bigr)=3Npt_{\textrm{2Q}}+\bigl[2(2N+1)p+1\bigr]t_{\textrm{1Q}}+t_{\textrm{mes}}+t_{\textrm{res}},$ (10)
[…]
@@
For $N=4,096$, the average size of light-cone-induced subgraphs is $N=5.67(2)$ at $p=1$, $N=16.613(5)$ at $p=2$, and $N=39.772(5)$ at $p=3$. Hence, in this regime, the average circuit execution time $t_{\textrm{circ}}$ is dominated by the passive reset time of the qubits. This is highlighted in Fig. 6E, where for a large number of samples, $t\simeq Kt_{\textrm{res}}$. For a small number of samples, the sampling time is instead dominated by the classical overhead. For a number of samples $K\gtrsim 10^{3}$, a first step for improving the sampling time is, therefore, to move to an active reset strategy for the qubits, which would yield $t_{\textrm{mes}}+t_{\textrm{res}}\simeq 6~\mu$[70], a $30$-fold reduction over passive reset.
@@
At the moment, for $K=10^{4}$, the total sampling time for a single light-cone-induced circuit takes about two to three seconds, as shown in Fig 6E. This is not competitive against classical state-vector emulations at $p=1$ and $p=2$, which can precondition all light-cone-induced circuits in less time (see Fig. 5). Thus, improvements will be needed to reduce the sampling time and make the approach competitive and capable of delivering an advantage when accounting for the preconditioning time budget. Whether this will prove a bottleneck remains to be established: indeed, we observe in Fig. 5 that the budget increases with the number of variables $N$ while the preconditioning is independent of $N$ for $p\ll\ln N$ thanks to caching of light-cone-induced subcircuits (see Appendix. G)—strictly speaking, preconditioning has a classical $O(N)$ dependence in finding the light-cone-induced subgraphs. If confirmed for much larger $N$, this scaling difference between the preconditioning budget and time could suggest a crossing value $N^{*}$ for a quantum simulation method (hardware or classically emulated) to deliver an advantage, independently of its intrinsic run-time.
@@ IV.2.2 Performance of a noisy quantum computer
where $F=f^{n_{\textrm{2Q}}}$ can be interpreted as the global quantum circuit fidelity, with individual errors happening independently and at random at the two-qubit gate level. To distinguish an $O(1)$ expectation value with $K$ sampled bit strings, one needs $K\sim F^{-2}$, or equivalently $K\sim f^{-3pN^{2}}$, according to the central limit theorem. By plugging in relevant numbers for $f$, NN, and $p$, one realizes that finite sampling by itself is likely not a bottleneck, but sampling for reliably estimating noisy and exponentially small expectation values might be. Therefore, hardware performance through $f$ sets the expectation for what circuit can be successfully executed. We verify that such a depolarizing noise model is able to phenomenologically fit experimental data in App. D, as similarly observed in Ref. [54].
@@ IV.2.3 Noisy quantum computer versus approximate classical tensor networks
Given the fidelity requirements to run the QAOA on large problem instances in the absence of quantum error correction, one will likely need to rely on a decomposition technique such as the light-cone one introduced in Sec. III.1.2. Light-cone-induced subgraphs at $p=3$ require at most $43$ qubits, far beyond what state-vector emulators can handle in a reasonable amount of time. Other candidate simulators for simulating such a quantum circuit include an actual quantum computer or classical approximate tensor-network methods, such as matrix product states [71, 72]. In the absence of quantum error correction, quantum computers are imperfect, and thus provide only an approximate simulation in some respects. A first question is whether quantum preconditioning via an approximately classically emulated preconditioner can deliver an advantage (see Appendix I).
Assuming an imperfect hardware output, a second question is what level of approximation in a tensor-network simulation matches the accuracy of the imperfect quantum computer? The level of approximation in tensor networks, encoded in a parameter $\chi\in[1,2^{N/2}]$ ,controls their run-time: the larger $\chi$, the less approximation is performed, and the longer the simulation takes. For instance, the algorithmic complexity of matrix product states goes as $O(\chi^{3})$. Then, one can answer the final question: for a matched output accuracy, which of the approximate tensor-network simulator or quantum computer is faster? This will provide a basis for a hardware-based quantum utility [73, 74, 75, 76].
@@
Many challenges faced by the proposed quantum-preconditioning method to deliver quantum utility are common to quantum optimization and are active areas of research. We discuss some of the main ones in the following.